What will be the angle at which the first dark band is formed in Fraunhofer diffraction patte de 0.3 mm wide and wavelength of source is 5890A
1
Expert's answer
2021-08-07T03:32:29-0400

"\sin\theta\approx \theta=\frac{\lambda}d,"

"\theta=0.0197~(rad)=1.12°."